# LISA: Layerwise Importance Sampling for Memory-Efficient Large Language Model Fine-Tuning

### v2

@@ Results
Table 2: The statistics of datasets. # Train and # Test denote the number of training and test samples respectively.
4.2 Moderate Scale Fine-Tuning
LISA is able to achieve this significant memory saving while still obtaining competitive performance in fine-tuning tasks.
Settings
To demonstrate the superiority of LISA over LoRA, we employ the instruction-following fine-tuning task with Alpaca GPT-4 dataset, which consists of 52k instances generated by GPT-4 (OpenAI et al., 2023) based on inputs from Alpaca (Taori et al., 2023). The effectiveness of fine-tuning was evaluated with MT-Bench (Zheng et al., 2023), featuring 80 high-quality, multi-turn questions designed to assess LLMs on multiple aspects.
In our experiments, we assessed five baseline models: GPT2-Small (Radford et al., 2019), TinyLlama (Zhang et al., 2024), Phi-2 (Li et al., 2023b), Mistral-7B (Jiang et al., 2023), LLaMA-2-7B, and LLaMA-2-70B (Touvron et al., 2023b). These models, varying in size with parameters ranging from 124M to 70B, were selected to provide a diverse representation of decoder models. Detailed specifications of each model are provided in Table 3. For hyperparameters, we adopt rank 128 for LoRA and E+H+2L for LISA in this section of experiments, with full details available in Appendix A.
Model Name Total Params Layers Model Dim Heads
GPT2-Small 124 M 12 768 12
@@ 4.2 Moderate Scale Fine-Tuning
Table 3: Baseline Model Specifications
Results
Table 4 offers a comprehensive evaluation of three fine-tuning methods—Full Parameter Fine-Tuning (FT), Low-Rank Adaptation (LoRA), and Layerwise Importance Sampling AdamW (LISA)—across a diverse set of tasks including Writing, Roleplay, Reasoning, Math, Extraction, STEM, and Humanities within the MT-Bench benchmark. The results clearly demonstrate LISA’s superior performance, which surpasses LoRA and full parameter tuning in most settings. Notably, LISA consistently outperforms LoRA and full parameter tuning in domains such as Writing, STEM, and Humanities. This implies that LISA can be beneficial for tasks involving memorization, while LoRA partially favors reasoning tasks.
MT-BENCH
Model & Method Writing Roleplay Reasoning Math Extraction STEM Humanities Avg. $\uparrow$
TinyLlama (Vanilla) 1.05 2.25 1.25 1.00 1.00 1.45 1.00 1.28
TinyLlama (FT) 3.27 3.95 1.35 1.33 1.73 2.69 2.35 2.38
TinyLlama (LoRA) 2.77 4.05 1.35 1.40 1.00 1.55 2.15 2.03
TinyLlama (LISA) 3.30 4.40 2.65 1.30 1.75 3.00 3.05 2.78
Mistral-7B (Vanilla) 5.25 3.20 4.50 2.70 6.50 6.17 4.65 4.71
Mistral-7B (FT) 5.50 4.45 5.45 3.25 5.78 4.75 5.45 4.94
Mistral-7B (LoRA) 5.30 4.40 4.65 3.30 5.50 5.55 4.30 4.71
Mistral-7B (LISA) 6.84 3.65 5.45 2.75 5.65 5.95 6.35 5.23
LLaMA-2-7B (Vanilla) 2.75 4.40 2.80 1.80 3.20 5.25 4.60 3.54
LLaMA-2-7B (FT) 5.55 6.45 3.60 2.00 4.70 6.45 7.50 5.18
LLaMA-2-7B (LoRA) 6.30 5.65 4.05 1.45 4.17 6.20 6.20 4.86
LLaMA-2-7B (LISA) 6.55 6.90 3.45 2.16 4.50 6.75 7.65 5.42
4.3 Large Scale Fine-Tuning
To further demonstrate LISA’s scalability on large-sized LLMs, we conduct additional experiments on LLaMA-2-70B (Touvron et al., 2023b).
MT-Bench GSM8K PubMedQA
Vanilla 5.69 54.8 83.0
FT 6.66 67.1 90.8
LoRA 6.52 59.4 90.8
LISA 7.05 61.1 91.6
Table 5: MT-Bench, GSM8K, and PubMedQA score for LLaMA-2-70B with three fine-tuning methods.
Settings
On top of the aforementioned instruction-following tasks in Section 4.2, we introduce an extra set of domain-specific fine-tuning tasks on mathematics and medical QA benchmarks. The GSM8K dataset (Cobbe et al., 2021), comprising 7473 training instances and 1319 test instances, was utilized for the mathematics domain. For the medical domain, we selected the PubMedQA dataset (Jin et al., 2019), which includes 211.3K artificially generated QA training instances and 1K test instances.
The statistics of these datasets are summarized in Table 2. Evaluation on the PubMedQA dataset (Jin et al., 2019) utilized a 5-shot prompt setting, while the GSM8K dataset (Cobbe et al., 2021) assessment was conducted using a Chain of Thought (CoT) prompt, as suggested by recent studies (Wei et al., 2022; Shum et al., 2023; Diao et al., 2023b). Regarding hyperparameters, we adopt rank 256 for LoRA and E+H+4L for LISA, where further details can be found in Appendix A.
Figure 5: Different aspects of LLaMA-2-70B model on MT-Bench.
Results
As shown in Table 5, LISA consistently produces better or on-par performance when compared with LoRA. Furthermore, in instruction-tuning tasks, LISA again surpasses full-parameter training, rendering it a competitive method for this setting. Specifically, Figure 5 highlights the model’s performance in various aspects, particularly LISA’s superiority over all methods in aspects of Writing, Roleplay, and STEM. On top of that, LISA displays consistently higher performance than LoRA on all subtasks, underscoring LISA’s effectiveness across diverse tasks. The chart also contrasts the yellow LoRA line with the purple Vanilla line, revealing that in large models like the 70B, LoRA does not perform as well as expected, showing only marginal improvements on specific aspects. More fine-tuning experiments result in the Appendix B.1.
4.4 Ablation Studies
Models MT-Bench Score Sampling Layers $\gamma$ Sampling Frequency $K$
TinyLlama 2.59 8 1
2.81 8 5
2.73 2 5
2.64 2 25
2.26 2 122
LLaMA-2-7B 4.94 8 1
5.11 8 5
4.91 2 5
4.88 2 25
4.64 2 122
Table 6: The MT-Bench scores derived from different LISA hyperparameters combination. All settings adopt learning rate $\eta_{0}=10^{-5}$.
Hyperparameters of LISA
The number of sampling layers $\gamma$ and sampling period $K$ are the two key hyperparameters of LISA. To obtain intuitive and empirical guidance of those hyperparameter choices, we conduct ablation studies using TinyLlama (Zhang et al., 2024) and LLaMA-2-7B (Touvron et al., 2023b) models with the Alpaca-GPT4 dataset. The configurations for $\gamma$, such as E+H+2L, E+H+8L, were denoted as $\gamma=2$ and $\gamma=8$. As for the sampling period — the number of updates per sampling interval — values were chosen from $\{1,5,25,122\}$, with 122 representing the maximum training step within our experimental framework. The findings, presented in Table 6, reveal that both $\gamma$ and $K$ markedly affect the LISA algorithm’s performance. Specifically, a higher $\gamma$ value increases the quantity of trainable parameters, albeit with higher memory costs. On the other hand, an optimal $K$ value facilitates more frequent layer switching, thereby improving performance to a certain threshold, beyond which the performance may deteriorate. Generally, the rule of thumb is: More sampling layers and higher sampling period lead to better performance. For a detailed examination of loss curves and MT-Bench results, refer to Appendix C.
Seed TinyLlama LLaMA-2-7B Mistral-7B
1 2.65 5.42 5.23
2 2.75 5.31 5.12
3 2.78 5.35 5.18
Sensitiveness of LISA
As LISA is algorithmically dependent on the sampling sequence of layers, it is intriguing to see how stable LISA’s performance is under the effect of randomness. For this purpose, we further investigate LISA’s performance variance over three distinct runs, each with a different random seed for layer selection. Here we adopt TinyLlama, LLaMA-2-7B, and Mistral-7B models with the Alpaca-GPT4 dataset, while keeping all other hyperparameters consistent with those used in the instruction following experiments in section 4.2. As shown in Table 7, LISA is quite resilient to different random seeds, where the performance gap across three runs is within $0.13$, a small value when compared with the performance gains over baseline methods.
5 Discussion
Theoretical Properties of LISA
Compared with LoRA, which introduces additional parameters and leads to changes in loss objectives, layerwise importance sampling methods enjoy nice convergence guarantees in the original loss. For layerwise importance sampled SGD, similar to gradient sparsification (Wangni et al., 2018), the convergence can still be guaranteed for unbiased estimation of gradients with increased variance. The convergence behavior can be further improved by reducing the variance with appropriately defined importance sampling strategy (Zhao and Zhang, 2015). For layerwise importance sampled Adam, there have been theoretical results in (Zhou et al., 2020) proving its convergence in convex objectives. If we denote $f$ as the loss function and assume that the stochastic gradients are bounded, then based on Loshchilov and Hutter (2017), we know that AdamW optimizing $f$ aligns with Adam optimizing $f$ with a scaled regularizer, which can be written as
where $\mathbf{S}$ is a finite positive semidefinite diagonal matrix. Following existing convergence results of RBC-Adam (Corollary 1 in Zhou et al. (2020)), we have the convergence guarantee of LISA in Theorem 1.
Theorem 1
where $f^{\rm reg}_{*}$ denotes the optimum value of $f^{\rm reg}$.
Better Importance Sampling Strategies
As suggested by the theoretical intuition, the strategy of E+H+2L in Section 4.2 and E+H+4L in Section 4.3 may not be the optimal importance sampling strategy, given it still sampled intermediate layers in a uniformly random fashion. We anticipate the optimizer’s efficiency to be further improved when taking data sources and model architecture into account in the importance sampling procedure.
6 Conclusion
In this paper, we propose Layerwise Importance Sampled AdamW (LISA), an optimization algorithm that randomly freezes layers of LLM based on a given probability. Inspired from observations of LoRA’s skewed weight norm distribution, a simple and memory-efficient freezing paradigm is introduced for LLM training, which achieves significant performance improvements over LoRA on downstream fine-tuning tasks with various models, including LLaMA-2-70B. Further experiments on domain-specific training also demonstrate its effectiveness, showing LISA’s huge potential as a promising alternative to LoRA for LLM training.
Limitations
The major bottleneck of LISA is the same as LoRA, where during optimization the forward pass still requires the model to be presented in the memory, leading to significant memory consumption. This limitation shall be compensated by approaches similar to QLoRA (Dettmers et al., 2023), where we intend to conduct further experiments to verify its performance.
Appendix A Training Setup and Hyperparameters
A.1 Training Setup
In our experiments, we employ the LMFlow toolkit (Diao et al., 2023a)***https://github.com/OptimalScale/LMFlow for conducting full parameter fine-tuning, LoRA tuning, and LISA tuning. We set the epoch number to 1 for both fine-tuning and continual pre-training scenarios. Additionally, we utilized DeepSpeed offload technology (Ren et al., 2021) to efficiently run the LLMs. All experiments were conducted on 8×8$\times$8 × NVIDIA Ampere Architecture GPU with 48 GB memory.
In our study, we explored a range of learning rates from $5\times 10^{-6}$ to $3\times 10^{-4}$, applying this spectrum to Full Parameter Training, LoRA, and LISA methods. For LoRA, we adjusted the rank $r$ to either 128 or 256 to vary the number of trainable parameters, applying LoRA across all linear layers. Regarding the number of sampling layers $\gamma$, our selections were guided by GPU memory considerations as reported in LoRA studies (Hu et al., 2022); For the LISA algorithm, we selected $\gamma=2$, and for experiments involving the 70B model, we opted for $\gamma=4$. The sampling period ($K$), defined as the number of update steps per sampling interval, ranged from 1 to 50. This range was influenced by variables such as the size of the dataset, the batch size, and the total number of training steps. To manage this effectively, we partitioned the entire training dataset into $K$ segments, thereby enabling precise regulation of the training steps within each sampling period.
A.2 Hyperparameter search
@@
Phi-2 $5\times 10^{-6}$ $5\times 10^{-5}$ 128 $5\times 10^{-5}$ 2 3
Mistral-7B $5\times 10^{-6}$ $5\times 10^{-5}$ 128 $5\times 10^{-5}$ 2 3
LLaMA-2-7B $5\times 10^{-6}$ $5\times 10^{-5}$ 128 $5\times 10^{-5}$ 2 3
LLaMA-2-70B $5\times 10^{-6}$ $5\times 10^{-5}$ 128 $5\times 10^{-5}$ 4 50
Table 8: The hyperparameter search identified optimal settings for each method: FP (Full Parameter Training), LoRA, and LISA.
We commenced our study with a grid search covering (i) learning rate, (ii) number of sampling layers $\gamma$, and (iii) sampling period $K$. Noting the effective performance of the LoRA method, we set the rank value to either $r=128$ or $r=256$.
The optimal learning rate was explored within the range $\{5\times 10^{-6},10^{-5},5\times 10^{-5},6\times 10^{-4},3\times 10^{-4}\}$, applicable to full parameter training, LoRA, and LISA.
@@
For the sampling period $K$, we examined values within ${1,3,5,10,50,80}$, aiming to maintain the model’s update steps within a range of 10 to 50 per sampling instance. This selection was informed by factors such as dataset size, batch size, and total training steps.
@@
Appendix B Additional Experimental Results
B.1 Instruction Following Fine-tuning
MT-BENCH
Model & Method Writing Roleplay Reasoning Math Extraction STEM Humanities Avg. $\uparrow$
LLaMA-2-70B(Vanilla) 7.77 5.52 2.95 1.70 6.40 7.42 8.07 5.69
LLaMA-2-70B(FT) 6.45 7.50 5.50 2.15 7.55 8.10 9.40 6.66
LLaMA-2-70B(LoRA) 7.55 7.00 5.30 2.60 6.55 8.00 8.70 6.52
LLaMA-2-70B(LISA) 8.18 7.90 5.45 2.75 7.45 8.60 9.05 7.05
Table 9 provides detailed MT-Bench scores for the LLaMA-2-70B model discussed in Section 4.3, demonstrating LISA’s superior performance over LoRA in all aspects under large-scale training scenarios. Furthermore, in Figure 6, we observe that LISA always exhibits on-par or faster convergence speed than LoRA across different models, which provides strong evidence for LISA’s superiority in practice.
It is also intriguing to observe from Figure 5 that Vanilla LLaMA-2-70B excelled in Writing, but full-parameter fine-tuning led to a decline in these areas, a phenomenon known as the “Alignment Tax” (Ouyang et al., 2022). This tax highlights the trade-offs between performance and human alignment in instruction tuning. LISA, however, maintains strong performance across various domains with a lower ”Alignment Tax“.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 6: Loss curves for LoRA, LISA and full-parameter tuning on the Alpaca-GPT4 dataset across different models.
Appendix C Ablation Experiments
C.1 Sampling Layers $\gamma$
We conducted an ablation study on the LLaMA-2-7B model trained with the Alpaca-GPT4 dataset, setting the sampling period $K=13$, so the number of samplings is exactly $10$. The study explored different configurations of sampling layers $\gamma$ including {E+H+2L, E+H+4L, E+H+8L}. Figure 7 depicts the impact of the number of sampling layers $\gamma$ on the training dynamics of the model. Three scenarios were analyzed: $\gamma=2$ (blue line), $\gamma=4$ (green line), and $\gamma=8$ (red line), throughout 120 training steps. Initially, all three configurations exhibit a steep decrease in loss, signaling rapid initial improvements in model performance. it’s clear that the scenario with $\gamma=8$ consistently maintains a lower loss compared to the $\gamma=2$ and $\gamma=4$ configurations, suggesting that a higher $\gamma$ value leads to better performance in this context. The radar graph of the MT-Bench score also indicates that the $\gamma 8$ configuration yields the best conversational ability of the model in this ablation experiment.
Figure 7: Comparison of loss curves for the $\gamma$ (number of sampling layers) ablation experiment.
C.2 Sampling Period $K$
Figure 8 displays the effects of varying sampling Period $K$ on training a 7B-sized model using the 52K-entry Alpaca-GPT4 dataset. This graph contrasts loss curves for different sampling period $K$ values: $K=122$ (green line), $K=25$ (red line), and $K=13$ (blue line) across 122 training steps. The results indicate that although each $K$ value results in distinct training trajectories, their convergence points are remarkably similar. This finding implies that for a 7B model trained on a dataset of 52K instruction conversation pairs, a sampling period of $K=13$ is optimal for achieving the best loss curve and corresponding MT-Bench score radar graph.
C.3 Sensitiveness to Randomness
LLaMA-2-7B on Alpaca-GPT4 with update step per sampling period $K=13$, and sampling layers $\gamma=2$, run 3 times with different random layer pick. Figure 9 shows that different random selections of layers slightly affect the training process but converge similarly. Despite initial fluctuations, the loss trends of three runs—distinguished by blue, green, and red lines—demonstrate that the model consistently reaches a stable state, underscoring the robustness of the training against the randomness in layer selection.
Appendix D Licenses
For instruction following and domain-specific fine-tuning tasks, all the datasets including Alpaca (Taori et al., 2023), GSM8k (Cobbe et al., 2021) and PubMedQA (Jin et al., 2019) are released under MIT license. For GPT-4, the generated dataset is only for research purposes, which shall not violate its terms of use.

### v3

@@ B.1 Training Setup
MMLU (Hendrycks et al., 2020) - 14,079
AGIEval (Zhong et al., 2023) - 9316
WinoGrande (Sakaguchi et al., 2021) - 44,000
@@
OpenWebMath (Paster et al., 2023) 6.3M -
Table 14: Baseline Model Specifications
Model Name # Params # Layers Model Dim # Heads
@@
Our study explored a range of learning rates from $5\times 10^{-6}$ to $3\times 10^{-4}$, applying this spectrum to Full Parameter Training, LoRA, and LISA methods. For LoRA, we adjusted the rank $r$ to either 128 or 256 to vary the number of trainable parameters, applying LoRA across all linear layers. Regarding the number of sampling layers $\gamma$, our selections were guided by GPU memory considerations as reported in LoRA studies (Hu et al., 2022); For the LISA algorithm, we selected $\gamma=2$, and for experiments involving the 70B model, we opted for $\gamma=4$. The sampling period ($K$), defined as the number of update steps per sampling interval, ranges from 1 to 50. This range was influenced by variables such as the size of the dataset, the batch size, and the number of training steps. To manage this effectively, we partitioned the entire training dataset into $K$ segments, thereby enabling precise regulation of the training steps within each sampling period.
B.2 Continual Pre-training Dataset
We extracted a high-quality subset from OpenWebMath (Paster et al., 2023), using the ‘Math_score’ attribute from the metadata as the metric for high-quality instances. The ‘Math_Score’ represents the probability that a document is mathematical, and we set the threshold at $0.95$. Finally, the number of tokens for this high-quality subset is 1.5 billion.
B.3 Hyperparameter search
We commenced our study with a grid search covering (i) learning rate, (ii) number of sampling layers $\gamma$, and (iii) sampling period $K$. Noting the effective performance of the LoRA method, we set the rank value to $r=128$ or $r=256$.
The optimal learning rate was explored within the range $\{5\times 10^{-6},10^{-5},5\times 10^{-5},6\times 10^{-4},3\times 10^{-4}\}$, applicable to full parameter training, LoRA, and LISA. For GaLore, we adhered to the official Transformers implementation***https://huggingface.co/blog/galore, utilizing default parameters, with the learning rate matching that of the full parameter training.
Table 15: The hyperparameter search identified optimal settings for each method: FP (Full Parameter Training), LoRA, GaLore, and LISA.
@@ B.3 Hyperparameter search
Mistral-7B $5\times 10^{-6}$ $5\times 10^{-5}$ 128 $5\times 10^{-5}$ 2 10
LLaMA-2-7B $5\times 10^{-6}$ $5\times 10^{-5}$ 128 $5\times 10^{-5}$ 2 10
LLaMA-2-70B $5\times 10^{-6}$ $5\times 10^{-5}$ 128 $5\times 10^{-5}$ 4 10
@@
For the sampling period $K$, we examined values within ${1,3,5,10,50,80}$, aiming to maintain the model’s update steps within a range of 10 to 50 per sampling period. This selection was informed by dataset size, batch size, and total training steps.
@@
Appendix C Licenses
For instruction following and domain-specific fine-tuning tasks, all the datasets, including Alpaca (Taori et al., 2023), GSM8k (Cobbe et al., 2021), MMLU (Hendrycks et al., 2020), AGIEval (Zhong et al., 2023) and PubMedQA (Jin et al., 2019) are released under MIT license. WinoGrande (Sakaguchi et al., 2021) and MT-Bench (Zheng et al., 2023) are under Apache-2.0 license. For GPT-4, the generated dataset is only for research purposes, which shall not violate its terms of use.